OpenStudy (anonymous):

Complete the solution of the equation. Find the value of y when x equals 13. -3x - 2y = -25 I get what to do, replace x with 13.. However -3 * 13 + -25.. Would that give me the answer? Or am I doing it wrong? @hartnn

hartnn (hartnn):

you mean - 3(13) +2y = -25 ?

OpenStudy (anonymous):

Oh lol. Sorry, now I see what I messed up on. Alright, so from there I would need to divide what 3(13) equals by 2, right?

hartnn (hartnn):

`divide what 3(13) equals by ` i didn't understand this -3(13) + 2y = -25 so, 2y = -25+3*13 got this ?

OpenStudy (anonymous):

Yes, I got that. Cause you're trying to solve for "Y" so you put y after the =. Correct?

hartnn (hartnn):

yeah all those steps to just isolate the y

OpenStudy (anonymous):

Ok, so that'd be the answer? "2y=-25+3*13"?

OpenStudy (anonymous):

Oh, wait.. Never mind. So to get rid of the 2 by the y, I divide by 2, on both sides?

hartnn (hartnn):

yes

hartnn (hartnn):

2y = 39-25 = 14 y = 14/2

OpenStudy (anonymous):

So 7 lol. Well, thanks :D that helped A LOT! You're literally a life saver :). Thanks again. I'll probably have more, I'll tag ya in em'.

hartnn (hartnn):

lol i won't get any notification :P

hartnn (hartnn):

but welcome ^_^

OpenStudy (anonymous):

Well, it turns out it was wrong? O_o

hartnn (hartnn):

because i took +2y instead of -2y

hartnn (hartnn):

so, y would be -7

OpenStudy (anonymous):

Awh darn lol. Well, thanks anyways :P.

hartnn (hartnn):

welcome ^_^